OpenStudy (3psilon):

Distance between two parallel lines 3x+y=12 and mx + 2y =4

OpenStudy (3psilon):

Wouldn't the slope of the second line have a slope of 3 also? It says the slope is 6 in my book

OpenStudy (anonymous):

Not 3 but -3

OpenStudy (amorfide):

parallel means it would have the same gradient perpendicular lines gradient would be the negative reciprocal

OpenStudy (amorfide):

yes it would be 3 also

OpenStudy (3psilon):

But why does it say "since they are parallel , m = 6" I'm confused

OpenStudy (anonymous):

Because you have a 2y in your equation.

OpenStudy (3psilon):

Wouldn't that make it y = 3/2x+2?

OpenStudy (3psilon):

Ohhh I see it now!

OpenStudy (radar):

Why wouldn't it be a -6 ??

OpenStudy (3psilon):

Yeah now that you mention it @radar I'm confused too 0_o

OpenStudy (radar):

And they want the perpendicular distance between these two parallel lines........right?

OpenStudy (3psilon):

Just the distance between?

OpenStudy (anonymous):

To start off, what do you need to obtain the distance?

OpenStudy (radar):

I believe the two equations can be resolved to: y=-3x + 12 and y=-3x + 2 these lines are parallel as they do have the same slope. But I don't know how to determine the distance between them. Maybe you determine a line that is perpendicular to one line and intersects at a point on the other. Good luck with it.

OpenStudy (3psilon):

D = Sqrt(10) @Chlorophyll ?

OpenStudy (anonymous):

Wow, that's fast! I haven't calculated at all, just express out my concept :P

OpenStudy (anonymous):

@3psilon However I'd like to see how do you come up with it, would you!

OpenStudy (3psilon):

Oh no it was in the book! lol

OpenStudy (anonymous):

You'd better roll up your sleeve now ;)

OpenStudy (3psilon):

How is 10/Sqrt(10) the same as sqrt(10)?

OpenStudy (anonymous):

Yep, my result is √10 units!

OpenStudy (anonymous):

Where do you get: 10/Sqrt(10) :0

OpenStudy (anonymous):

Anyway, 10/√10 = √10 * √10 / √10 = ....

OpenStudy (radar):